OpenStudy (anonymous):

A substance of unknown mass absorbs 1932 joules of energy, going from 26 to 42 degrees Celsius. If the specific heat of the substance is 4.29 J/g·°C, how much of the substance is present?

OpenStudy (anonymous):

DH= m cp DT = 4,29 J/gC * 16 °C = 68,64 *m = 1932 J -> m= 1932 J /68,64 J/g = 28,14 g